# Features of the EAR2 neutron beam following the spallation target upgrade at the n_TOF facility at CERN

### v1

@@ Abstract
The n_TOF facility at CERN has undergone a major upgrade after the installation of a new spallation target, designed to improve the features of both neutron beamlines at the experimental areas 1 and 2 (EAR1 and EAR2) and the commissioning of a new experimental area (NEAR). Due to improved coupling of the spallation target with the EAR2 beamline, the upgrade resulted in a significantly increased neutron flux and improved neutron energy resolution. This paper presents the results of the commissioning phase that followed to characterise the EAR2 neutron beamline and validate the FLUKA Monte Carlo simulations of the facility. The main features of the neutron beam, namely the neutron flux, spatial profile and energy resolution, are evaluated and compared to the previous target. The neutron flux presents a general increase of 20% below 1 eV, 40% between 1 eV and 100 keV and 50% between 100 keV and 10 MeV. The measured width of the beam profile was 3 cm (FWHM) at the reference position for neutron capture measurements. The energy resolution with the new spallation target shows a significant improvement compared to the previous one. Moreover, FLUKA Monte Carlo simulations present a good agreement with the measured neutron flux and profile within uncertainties, and a remarkable reproduction of the energy resolution.
@@ 1 Introduction
The nTOF Collaboration operates the neutron time-of-flight (TOF) facility at CERN, based on a 20 GeV/c pulsed proton beam impinging on a solid lead target where water is employed to moderate the neutrons produced by spallation reactions. The facility is characterized by a high-instantaneous neutron beam intensity, high energy resolution, and a wide neutron energy spectrum, spanning from sub-thermal to GeV. The scientific activities of the n_TOF Collaboration are mostly focused on the measurement of neutron-induced cross sections of interest in astrophysics [1], nuclear technology [2] and medical physics [3]. The first experimental area, EAR1, in operation since 2001, is located at 185 m from the spallation target nearly in the same direction as the incoming proton beam. In 2014, a new experimental area, EAR2, located 20 m above the target perpendicular to the proton beam direction, was commissioned.
During the CERN Long Shutdown (2019-2021), the facility has gone through a major upgrade including the installation of a new spallation target designed to fully optimise the features of the nTOF experimental areas, unlike the previous one specifically designed for EAR1 [4]. Moreover, the development of a new experimental area (NEAR) has been completed. NEAR [5] is located at 3 m to the left of the target with respect to the proton beam direction, aiming to explore neutron reactions of interest in astrophysics and radiation damage induced in materials by neutrons.
The new target is expected to improve the features of EAR2 neutron beamline in terms of neutron flux and energy resolution. However, a good knowledge of every characteristic of the neutron beam is necessary for the correct analysis of the experimental data. For this purpose, an extensive commissioning campaign for the neutron beam characterization has been carried out.
@@ 2 EAR2: the n_TOF 20 m neutron beamline
EAR2 is located above the spallation target at about 20 m along the vertical direction. Figure 1 shows a drawing of the beamline’s main components. From bottom to top: solid lead spallation target, water moderator, vacuum line window, first collimator, sweeping magnet, filter station, second collimator with two configurations depending on the experimental requirements (small and big, of 21.8 and 60 mm downstream inner diameter, respectively), the experimental area with a lead collimator (only in combination with the small collimator) and the beam dump. The second collimator has a conical shape, hence why the width of the beam profile widens as it moves away from it (in more detail in Section 5).
The EAR2 beamline was completed in 2014 [6, 7], and it was in operation for 5 years using the existing spallation target that was optimised only for EAR1. During this phase, several measurements were performed [8, 9, 10, 11, 12] exploiting the advantage of a high instantaneous flux, i.e., high number of neutrons per pulse, that allows measuring cross sections of samples of very low masses compared to EAR1, as the beam time required for the measurement is lower and, most importantly, as the signal-to-background ratio for very radioactive samples is much higher.
@@ 3 Monte Carlo simulations with FLUKA
Monte Carlo (MC) simulations allow to assess neutron beam characteristics such as the neutron flux and the spatial beam profile and are essential to determine the energy resolution of a TOF instrument over a wide range of neutron energies. During the commissioning phase, the simulations are validated against experimental measurements and hence become a powerful tool to further improve and optimise the quality of the neutron beam and other features of the experimental area such as the backgrounds. Moreover, they are indispensable in the accurate determination of the relation between energy and TOF, as discussed in detail in Sections 4 and 6. In this section, we present the implementation and technical details of the MC simulations.
@@
The simulation of the whole spallation process induced by protons impinging on the lead target is required to track and record the position, direction, energy and time-of-flight of every neutron arriving to the experimental hall. The FLUKA code [14, 15, 16] version 4.3 has been used to carry out these simulations. This version includes a point wise treatment of the neutron cross sections below 20 MeV [17], while neutron interactions above this energy are managed by FLUKA’s internal nuclear models, allowing a precise production and transport of the neutrons in the whole energy range of interest for n_TOF. A detailed implementation of the geometry of the whole facility, based on its blueprints, has been carried out since the first phase of operation. For the new spallation target, the geometry of the previous phase has been adapted and updated accordingly, including a detailed description of all the constituent materials. In Figure 3 we show a visualisation of the geometry implemented in FLUKA using its graphical user interface Flair [18].
@@
In terms of CPU time, these simulations are time-consuming and a full MC transport through the entire beamline is not feasible at the moment. To improve the speed of the simulations without compromising the reliability of the results, we have considered a two steps simulation approach. The positions, direction, energy and time relative to the generation of the neutrons and -rays emitted from the target in the spallation process are scored on the hemispherical surface of the vacuum pipe window, at the interface from the aluminium window to vacuum, that is indicated with a blue arrow in Figure 3. Within a sufficiently small angle ( for EAR2 [19]), relative to the neutron beam pipe axis, the neutron’s angular distribution can be assumed to be isotropic. We continue the neutron transport up to the experimental hall by optically propagating the neutrons along the beamline, performing a sampling of the incident angle within . This optical transport is carried out by a C++ custom code [19, 20]. The collimation system is modelled in such a way that, if a neutron hits the collimator, it is discarded and does not arrive in the experimental hall. In this way, we can also reproduce the spatial distribution of the neutron beam as shown in Section 5.
@@ 4 Neutron flux
At n_TOF we do not define neutron flux, , in a traditional way as the number of neutrons per unit of time/second but per nominal pulse, which is produced by protons. Therefore,, is defined as the number of neutrons per neutron energy, , integrated over the entire spatial profile of the neutron beam arriving to the experimental area and expressed as per nominal proton pulse. Here refers to the neutron intensity, but for simplicity we will refer to it hereafter as the “neutron flux”. To reach a high accuracy of time-of-flight technique, a precise knowledge of the neutron flux at the sample position is required.
To evaluate the neutron flux at n_TOF-EAR2, we have carried out a commissioning campaign combining many complementary detection systems and different standard cross sections. In this section, we present the experimental setup and the methodology adopted. Moreover, we include a comparison with simulations and with the flux available with the previous target.
@@ 4.1 Methodology
where ) is the number of counts registered, ) is the background contribution, and is the total detection efficiency. Combining equations (1) and (2), it is possible to extract the neutron flux:
@@
where is the neutron rest mass and is an effective neutron path before reaching the hemispherical vacuum window, that accounts for the stochastic nature of the moderation process [20]. In fact, is a stochastic quantity with a distribution shown in Fig 4. It means that a distribution of corresponds to a specific measured , or conversely, neutrons with the same arrive at a different . The relation of and (or ) is then called energy resolution function of the facility and is discussed in detail in Sec. 6.
@@
According to Eq. (3), the neutron flux is extracted by dividing the number of counts (per proton pulse) by the theoretical yield. In the latter, any involved cross section is convoluted with the distribution, obtained via MC simulations of the spallation process carried out with FLUKA, in order to account for the effect of the RF. The flight path in Eq. (5) is determined by means of an iterative process where the position of the resonances in the in time is compared to the ones in the theoretical yield convoluted with the RF. In the absence of resonances, the dips caused by the neutron absorption in the aluminium windows of the target can serve for the same purpose. The result of this process—when the resonances (or dips) match in both yields—gives an effective flight path, , that accounts for the effect of the RF. Typically, the same flight path can be used for the whole measured energy. However, there are cases in which multiple flight paths need to be used for different energy ranges.
4.2 Evaluation
@@
The experimental setup consisted of solid state detectors, SiMon2 (Silicon Monitor) [21], and gaseous detectors, MGAS (Micro-mesh Gaseous Structure) [22, 23, 24] and PPACMon (Parallel Plate Avalanche Counters) [25, 26]. The use of different detector systems aims to reduce the systematic uncertainties. The commissioning measurements were carried out with the small collimator, since it is the most frequently used one.
@@
Figure 5 shows the setup in the experimental hall. The neutron beam comes from the bottom, passing through SiMon2, MGAS and PPACMon before being stopped at the beam dump downstream (not shown), placed above the ceiling. The isotopes, reactions and sample masses used for every detector, together with the ranges of energy in which they have been used are summarised in Table 1. The sizes of all the samples employed in the flux evaluation are large enough to cover the whole neutron beam profile. The energy region from 25 meV to 200 keV is covered by SiMon2 with 6Li and from 25 meV to 30 keV by MGAS with 10B. The 6Li(n,t) and 10B(n,) capture cross sections are considered standard between 25 meV and 1 MeV [27], although due to the effects of the -flash and pile-up in our measurements, the first is limited to 200 keV and the latter to 30 keV. The values of the flux below 25 meV presented are also obtained by a combination of these two datasets, since there were no other standards available. The 235U(n,f) cross section is considered standard within the energy range of 0.15 - 200 MeV; however, we have decided to extend the use of 235U data down to 30 keV, since the 235U(n,f) cross section is smooth and well-known in this energy range. MGAS 235U is reliable up to 5 MeV because of the -flash, and from 1 MeV it uses PPACMon 235U data as a TOF reference due to the impact of the slower MGAS signals on the initial time determination in such energies. The energy range from 5 MeV up to 200 MeV is then covered only with PPACMon 235U and 238U, while above 200 MeV the count rate is very low for achieving reliable statistics. Furthermore, since the 235U sample of PPACMon has the best-characterised mass, with an accuracy of only 0.4%, it has been used for absolute normalization of the flux. In particular, the flux measured with all the detectors has been normalised to the integral of the PPACMon flux in the 7.8-11 eV range, which for 235U(n,f) is also considered a standard [27].
@@
Figure 7 shows the statistical and systematic uncertainties as a function of the neutron energy . The systematic uncertainty for each of the measurements has been calculated by means of error propagation, considering the uncertainty of the different corrections applied during the data analysis, as well as the uncertainties related to the use of evaluated cross sections and MC simulations. Above 10 MeV the total number of neutrons arriving into the experimental hall decreases considerably. Therefore, the difficulty of detecting neutrons is bigger and the statistical error dominates the uncertainty. The jump in statistical uncertainties at 10 MeV is due to the use of a single detector in the evaluation in this range. The same effect can be seen in the systematic uncertainties. As we go down in energy, the systematic uncertainties are more relevant. Below 10 keV the systematic uncertainties come mainly from the MC simulations used in the calculation of the flux. This uncertainty is the same in the whole energy range. The flux derived from different detectors is also fully consistent within the statistical uncertainties. Table 2 shows the integral value of the neutron flux per proton pulse in each decade of the energy spectrum and the average of the systematic and statistical uncertainties in the same decade.
@@
Figure 8, top panel shows the evaluated neutron flux along with the results of the FLUKA simulations carried out in two steps (production of spallation neutrons, and resampling and propagation up to the experimental hall). The bottom panel then shows the ratio between FLUKA and the evaluated flux. The absolute value of the simulated flux is about 20% higher with respect to the experimental data, similar ratio as obtained with the previous spallation target [29]. The shape of the neutron energy flux is in general well reproduced by the simulations, which after scaling to the evaluation, are in agreement within a 3% in the range from 1 eV to 100 keV, while in the regions below 1 eV and from 100 keV up to 20 MeV the agreement is within a 10%. These differences might arise from possible imperfections in the modeling of the geometry and materials, as well as small misalignments in elements in the actual beamline such as collimators. Above 20 MeV the simulations underestimate the experimental flux, an effect that we also observe in EAR1 and NEAR. In this energy range the simulations rely only on theoretical models of the neutron generation and transport, and the origin of the inconsistencies is unclear.

### v2

# Characterisation of the neutron beam in the n_TOF-EAR2 experimental area at CERN following the spallation target upgrade

@@ Abstract
The n_TOF facility at CERN has undergone a major upgrade after the installation of a new spallation target, designed to improve the performance of both neutron beamlines at the experimental areas 1 and 2 (EAR1 and EAR2) and the commissioning of a new experimental area (NEAR). Due to improved coupling of the spallation target with the EAR2 beamline, the upgrade resulted in a significantly increased neutron flux and improved neutron energy resolution. This paper presents the results of the commissioning phase that followed to characterise the EAR2 neutron beamline and validate the FLUKA Monte Carlo simulations of the facility. The main characteristics of the neutron beam, namely the neutron flux, spatial profile and energy resolution, are evaluated and compared to the previous target. The neutron flux presents a general increase of 20% below 1 eV, 40% between 1 eV and 100 keV and 50% between 100 keV and 10 MeV. The measured width of the beam profile was 3 cm full width at half maximum (FWHM) at the reference position for neutron capture measurements. The energy resolution with the new spallation target shows a significant improvement compared to the previous one. Moreover, FLUKA Monte Carlo simulations present a good agreement with the measured neutron flux and profile within uncertainties, and a remarkable reproduction of the energy resolution.
@@ 1 Introduction
The nTOF Collaboration operates the neutron time-of-flight (TOF) facility at CERN, based on a 20 GeV/c pulsed proton beam impinging on a solid lead target where water is employed to moderate the neutrons produced by spallation reactions. The facility is characterized by a high-instantaneous neutron beam intensity, high energy resolution, and a wide neutron energy spectrum, spanning from sub-thermal to GeV. The scientific activities of the n_TOF Collaboration are mostly focused on the measurement of neutron-induced cross sections of interest in astrophysics [1], nuclear technology [2] and medical physics [3]. The first experimental area, EAR1, in operation since 2001, is located at 185 m from the spallation target nearly in the same direction as the incoming proton beam. In 2014, a second experimental area was commissioned, EAR2, located 20 m above the target perpendicular to the proton beam direction.
During the CERN Long Shutdown (2019-2021), the facility has gone through a major upgrade including the installation of a new spallation target designed to fully optimise the capabilities of the nTOF experimental areas, unlike the previous one specifically designed for EAR1 [4]. Moreover, the development of new experimental area, NEAR, has been completed. NEAR [5] is located at 3 m to the left of the target with respect to the proton beam direction, aiming to explore neutron reactions of interest in astrophysics and radiation damage induced in materials by neutrons.
The new target is expected to improve the performance of the EAR2 neutron beamline in terms of neutron flux and energy resolution. However, a good knowledge of every characteristic of the neutron beam is necessary for the correct analysis of the experimental data. For this purpose, an extensive commissioning of the neutron beam characterization has been carried out.
@@ 2 EAR2: the n_TOF 20 m neutron beamline
EAR2 is located above the spallation target at about 20 m along the vertical direction. Figure 1 shows a drawing of the beamline’s main components. From bottom to top: solid lead spallation target, water moderator, vacuum line window, first collimator, sweeping magnet, filter station, second collimator with two configurations depending on the experimental requirements (small and big, of 21.8 and 60 mm downstream inner diameter, respectively), the experimental area with a lead collimator (only in combination with the small collimator) and the beam dump. The second collimator has a conical shape, hence why the width of the beam profile widens after exiting the collimator (in more detail in Section 5). The first collimator is made of steel, while the second collimator is made of steel and polyethylene (including a boron carbide insert in the small configuration).
The EAR2 beamline was completed in 2014 [6, 7], and it was in operation for 5 years using the existing spallation target that was optimised only for EAR1. During this phase, several measurements were performed [8, 9, 10, 11, 12] exploiting the advantage of a high instantaneous flux, i.e., high number of neutrons per pulse, that allows measuring cross sections of samples of very low masses compared to EAR1, as the beam time required for the measurement is shorter and, most importantly, as the signal-to-background ratio for very radioactive samples is much higher. A comparison of pulsed neutron sources was carried out by Colonna et al. [13], where EAR2 showed the highest neutron flux per proton pulse, with respect to the main facilities in operation worldwide.
@@ 3 Monte Carlo simulations with FLUKA
Monte Carlo (MC) simulations allow to assess neutron beam characteristics such as the neutron flux and the spatial beam profile and are essential to determine the energy resolution of a TOF instrument over a wide range of neutron energies. During the commissioning phase, the simulations are validated against experimental measurements, thereby becoming a powerful tool to further improve and optimise the quality of the neutron beam as well as other aspects of the experimental area such as the background conditions. Moreover, they are indispensable in the accurate determination of the relation between energy and TOF, as discussed in detail in Sections 4 and 6. In this section, we present the implementation and technical details of the MC simulations.
@@
The simulation of the whole spallation process induced by protons impinging on the lead target is required to track and record the position, direction, energy and TOF of every neutron generated in the MC simulation and arriving in the experimental hall. The FLUKA code [15, 16, 17] version 4.3 has been used to carry out these simulations. This version includes a point wise treatment of the neutron cross sections below 20 MeV [18], while neutron interactions above this energy are managed by FLUKA’s internal nuclear models, allowing a precise production and transport of the neutrons in the whole energy range of interest for n_TOF. A detailed implementation of the geometry of the whole facility, based on its blueprints, has been carried out since the first phase of operation. For the new spallation target, the geometry of the previous phase has been adapted and updated accordingly, including a detailed description of all the constituent materials. In Figure 3 we show a visualisation of the geometry implemented in FLUKA using its graphical user interface Flair [19].
@@
In terms of CPU time, these simulations are time-consuming and a full MC transport through the entire beamline is not feasible at the moment. To improve the speed of the simulations without compromising the reliability of the results, we have considered a two steps simulation approach. The positions, direction, energy and time relative to the generation of the neutrons and -rays emitted from the target in the spallation process are scored on the hemispherical surface of the vacuum pipe window, at the interface from the aluminium window to vacuum, that is indicated with a blue arrow in Figure 3. Within a sufficiently small angle ( for EAR2 [20]), relative to the neutron beam pipe axis, the neutron’s angular distribution can be assumed to be isotropic. We continue the neutron transport up to the experimental hall by optically propagating the neutrons along the beamline, performing a sampling of the incident angle within . This optical transport is carried out by a C++ custom code [20, 21]. The collimation system is modelled in such a way that, if a neutron hits the collimator, it is discarded and does not arrive in the experimental hall. In this way, we can also reproduce the spatial distribution of the neutron beam as shown in Section 5. The neutrons arriving to the experimental area at the requested flight path are computed according to their position, energy and TOF to obtain the quantities of interest.
@@ 4 Neutron flux
The definition of flux, , according to the International Commission of Radiation Units and Measurements (ICRU) [22] is
where is the increment of the number of neutrons with energy between and in the time interval , with units of .
At n_TOF the measurements are normalised per total number of protons on target rather than per duration of the measurement in seconds, since the proton pulse frequency is not constant and can differ considerably from the average 0.8 Hz. Therefore, the definition of flux used at n_TOF is the number of neutrons with energy between and , integrated over the entire profile of the beam arriving to the experimental area and per proton on target,
with units of and scaled to the nominal proton pulse, which is produced by protons.
A precise knowledge of the neutron flux at the sample position is required for achieving highly accurate measurements. For this reason, we have carried out a commissioning campaign measurement of the flux combining several complementary detection systems and different standard cross sections. We refer to flux evaluation to this way of determining the flux, and evaluated flux to the high precision measured flux, in order to differentiate from the often less precise flux monitored online during the measurements.
@@ 4.1 Methodology
where ) is the number of counts registered, ) is the background contribution, and is the total detection efficiency, all of them expressed per energy unit and integrated per nominal proton pulse. Combining equations (3) and (4), it is possible to extract the neutron flux:
@@
where is the neutron rest mass and is an effective neutron path before reaching the hemispherical vacuum window, that accounts for the stochastic nature of the moderation process [21]. In fact, is a stochastic quantity with a distribution shown in Fig 4. It means that a distribution of corresponds to a specific measured , or conversely, neutrons with the same arrive at a different . The relation of and (or ) is then called energy resolution function (RF) of the facility and is discussed in detail in Sec. 6.
@@
According to Eq. (5), the neutron flux is extracted by dividing the number of counts (per proton pulse) by the theoretical yield. In the latter, any involved cross section is convoluted with the distribution, obtained via MC simulations of the spallation process carried out with FLUKA, in order to account for the effect of the RF. The flight path in Eq. (7) is determined by means of an iterative process where the position of the resonances in the in time is compared to the ones in the theoretical yield convoluted with the RF. In the absence of resonances, the dips caused by the neutron absorption in the aluminium windows of the target can serve for the same purpose. The result of this process—when the resonances (or dips) match in both yields—gives an effective flight path, that accounts for the effect of the RF. Typically, the same flight path can be used for the whole measured energy. However, there are cases in which multiple flight paths need to be used for different energy ranges.
4.2 Flux evaluation
@@
The experimental setup consisted of solid state detectors, SiMon2 (Silicon Monitor) [23], and gaseous detectors, MGAS (Micro-mesh Gaseous Structure) [24, 25, 26] and PPACMon (Parallel Plate Avalanche Counters) [27, 28]. The details on the analysis of the different reactions is specified in the detector references. The use of different detector systems aims to reduce the systematic uncertainties. The commissioning measurements were carried out with the small collimator, since it is the most frequently used one.
@@
Figure 5 shows the setup in the experimental hall. The neutron beam comes from the bottom, passing through SiMon2, MGAS and PPACMon before being stopped at the beam dump downstream (not shown), placed above the ceiling. The isotopes, reactions and sample masses used for every detector, together with the ranges of energy in which they have been used are summarised in Table 1. The sizes of all the samples employed in the flux evaluation are large enough to cover the whole neutron beam profile. The energy region from 25 meV to 200 keV is covered by SiMon2 with 6Li and from 25 meV to 30 keV by MGAS with 10B. The 6Li(n,t) and 10B(n,) capture cross sections are considered standard between 25 meV and 1 MeV [29], although due to the effects of the -flash and pile-up in our measurements, the first is limited to 200 keV and the latter to 30 keV. In particular, in the case of 6Li(n,t), the effect of the angular distribution of the reaction above 100 keV was taken into account in the estimation of the detector efficiency via simulations. The values of the flux below 25 meV presented are also obtained by a combination of these two datasets, since there were no other standards available. The 235U(n,f) cross section is considered standard within the energy range of 0.15 - 200 MeV; however, we have decided to extend the use of 235U data down to 30 keV, since the 235U(n,f) cross section is smooth and well-known in this energy range. MGAS 235U is reliable up to 5 MeV because of the -flash, and from 1 MeV it uses PPACMon 235U data as a TOF reference due to the impact of the slower MGAS signals on the initial time determination in such energies. The energy range from 5 MeV up to 200 MeV is then covered only with PPACMon 235U and 238U, while above 200 MeV the count rate is very low for achieving reliable statistics. Furthermore, since the 235U sample of PPACMon has the best-characterised mass, with an accuracy of only 0.4%, it has been used for absolute normalization of the flux. In particular, the flux measured with all the detectors has been normalised to the integral of the PPACMon flux in the 7.8-11 eV range, which for 235U(n,f) is also considered a standard [29].
@@
Figure 7 shows the statistical and systematic uncertainties as a function of the neutron energy . The systematic uncertainty for each of the measurements has been calculated by means of error propagation, considering the uncertainty of the different corrections applied during the data analysis, as well as the uncertainties related to the use of evaluated cross sections and MC simulations. Above 10 MeV the total number of neutrons arriving into the experimental hall decreases considerably. Therefore, the difficulty of detecting neutrons is bigger and the statistical error dominates the uncertainty. The jump in statistical uncertainties at 10 MeV is due to the use of a single detector in the evaluation in this range. The same effect can be seen in the systematic uncertainties. As we go down in energy, the systematic uncertainties are more relevant. Below 10 keV the systematic uncertainties come mainly from the MC simulations used in the calculation of the flux. This uncertainty is the same in the whole energy range. The flux derived from different detectors is also fully consistent within the statistical uncertainties. The evaluated flux presents a series of dips between 50 and 200 keV, due to neutron transmission through the aluminium windows in the target-moderator assembly and the beam line, which present resonances in its cross sections. The effect of these dips is minimal for the measurement of actinide cross sections in their unresolved resonance region, while for measurements of cross sections with resonances this energy range is already constrained by the RF. Table 2 shows the integral value of the neutron flux per proton pulse in each decade of the energy spectrum and the average of the systematic and statistical uncertainties in the same decade.
@@
Figure 8, top panel shows the evaluated neutron flux along with the results of the FLUKA simulations carried out in two steps (production of spallation neutrons, and resampling and propagation up to the experimental hall). The bottom panel then shows the ratio between FLUKA and the evaluated flux. The absolute value of the simulated flux is about 20% higher with respect to the experimental data, similar ratio as obtained with the previous spallation target [31]. The shape of the neutron energy flux is in general well reproduced by the simulations, which after scaling to the evaluation, are on average in agreement within a 3% per energy decade in the range from 1 eV to 100 keV, while in the regions below 1 eV and from 100 keV up to 20 MeV the average agreement is within a 10%. These differences might arise from possible imperfections in the modeling of the geometry and materials, as well as small misalignments in elements in the actual beamline such as collimators. Above 20 MeV the simulations underestimate the experimental flux, an effect that we also observe in EAR1 and NEAR. In this energy range the simulations rely only on theoretical models of the neutron generation and transport, and the origin of the inconsistencies is unclear.